Determine the complete factorization for the function g(x)=x^4 - 16
I believe that it is (x^2+8) (x^2-2) but I’m not really sure I’m how to do this. Please help me answer this I learned it about three years ago but I forgot how to.

Answer by ikleyn(52780)

.

x^4 - 16 = (x^2-4)*(x^2+4) = (x-2)*(x+2)*(x^2+4)


is the most complete possible factoring over real numbers.

Solved,  answered and explained.